OpenStudy (anonymous):

Can someone help me with this. I dont know how to do this im not good at math. The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today? 900 620 520 450

OpenStudy (anonymous):

idk fool

OpenStudy (zale101):

a + b = 1070 3a - 2b = 110

OpenStudy (zale101):

2 * (a + b) + 3a - 2b = 2 * 1070 + 110

OpenStudy (zale101):

can u continue from up here?

OpenStudy (anonymous):

hmmm. is it 2a + 2b = 1,180

OpenStudy (zale101):

2a + 2b + 3a - 2b = 2140 + 110

OpenStudy (anonymous):

2a + 2b + 3a - 2b = 2240 ?

OpenStudy (zale101):

no, it's 5a = 2250

OpenStudy (anonymous):

448?

OpenStudy (anonymous):

no 450

OpenStudy (zale101):

yes

OpenStudy (anonymous):

thank you imma post another can you help me ?